OpenStudy (anonymous):

> Double checking my answer < Write the standard form of the equation y+4=-12/7(x-1). 12x+7y=-16

OpenStudy (anonymous):

does the first equation look like \[y+4=\frac{ -12 }{ 7(x-1) }\]?

OpenStudy (anonymous):

-12/7 is a fraction.

OpenStudy (anonymous):

so \[y+4=\frac{ -12 }{ 7}\left( x-1 \right)\]

OpenStudy (anonymous):

Ya.

OpenStudy (anonymous):

yes you are correct

OpenStudy (anonymous):

could also be written as: \[y=-\frac{ 12}{ 7 }x-\frac{ 16 }{ 7 }\] or \[y=\frac{ -12x-16 }{ 7 }\]